Calculate [H3O+] given [OH-] in each aqueous solution

part a ) [OH-] = 5.7 x 10^-12 M and classify if acidic or basic

part b) [OH-]= 2.0 x 10^-2 M and classify if acidic or basic

part c) [OH-] = 6.1 x 10^-10 M and classify if acidic or basic

part d) [OH-] = 4.3 x 10^-4 M and classify if acidic or basic
Answer #1

we know in any solution [H3O+][OH-] = Kw

A solution is acidic if [H3O+] > 1.0x10-7 M

Calculate [H3O+] given [OH-] in each aqueous solution

part a ) [OH-] = 5.7 x 10^-12 M and classify if acidic or basic

[H3O+] = Kw / [OH-] as we know in any solution [H3O+][OH-] = Kw

So

[H3O+] = 1.0x10-14 / 5.7x10-12 = 1.75 x10-3 M and the solution is acidic

part b) [OH-]= 2.0 x 10^-2 M and classify if acidic or basic

[H3O+] = Kw / [OH-] = 1.0x10-14 / 2.0x10-2 = 5.0x10-13 M and the solution is Basic

part c) [OH-] = 6.1 x 10^-10 M and classify if acidic or basic

[H3O+] = Kw / [OH-] = 1.0x10-14 / 6.1x10-10 =  1.639x10-5M and the solution is acidic

part d) [OH-] = 4.3 x 10^-4 M and classify if acidic or basic

[H3O+] = Kw / [OH-] = 1.0x10-14 / 4.3x10-4 = 2.325x10-11 M and the solution is basic